# Multi-fidelity learning for interatomic potentials: Low-level forces and high-level energies are all you need

### v2

@@ Acknowledgments
All models in this study utilize the “Hierarchically Interacting Particle Neural Network with Tensor Sensitivity Information” (HIP-NN-TS) architecture, [38] an extension of the original HIP-NN architecture. [52] The original HIP-NN architecture is limited in terms of how information about atomic neighborhoods is collected by the model. The extended HIP-NN-TS architecture can incorporate either scalar, dipole, or quadrupole information about the feature distribution in the environment of an atom. HIP-NN-TS with tensor sensitivity of zero (i.e., scalar information) is equivalent to the original HIP-NN model. The accuracy (and computational cost) of HIP-NN-TS increases somewhat with increasing tensor sensitivity. To maximize accuracy, all HIP-NN-TS models reported in the main text utilize tensor sensitivity of $\ell=2$ (that is, using scalar, dipole, and quadrupole information). For comparison, Supporting Information provides results for other HIP-NN-TS models with tensor sensitivity of zero (scalar information) and one (scalar and dipole information). For brevity, we refer to HIP-NN-TS simply as HIP-NN throughout the main text.

### v3

@@ 5 Conclusions and Future Work
Data availability
The data that support the findings of this study will be openly available following an embargo at the following URL/DOI: https://github.com/lanl/hippynn/examples/multifidelity_ani1x_training.py.
@@ Acknowledgments
Not all multi-fidelity learning approaches utilize multi-head learning. For example, an alternative multi-fidelity approach is to provide the desired level of theory directly to the neural network as an input, such that a single output head predicts the energy corresponding to the specified level of theory. [89]
@@
All models in this study utilize the “Hierarchically Interacting Particle Neural Network with Tensor Sensitivity Information” (HIP-NN-TS) architecture, [38] an extension of the original HIP-NN architecture. [53] The original HIP-NN architecture is limited in terms of how information about atomic neighborhoods is collected by the model. The extended HIP-NN-TS architecture can incorporate either scalar, dipole, or quadrupole information about the feature distribution in the environment of an atom. HIP-NN-TS with tensor sensitivity of zero (i.e., scalar information) is equivalent to the original HIP-NN model. The accuracy (and computational cost) of HIP-NN-TS increases somewhat with increasing tensor sensitivity. To maximize accuracy, all HIP-NN-TS models reported in the main text utilize tensor sensitivity of $\ell=2$ (that is, using scalar, dipole, and quadrupole information). For comparison, Supplementary Material provides results for other HIP-NN-TS models with tensor sensitivity of zero (scalar information) and one (scalar and dipole information). For brevity, we refer to HIP-NN-TS simply as HIP-NN throughout the main text.